Find the value of each of the following quantities:

C(6,1)=

C(8,1)= 

C(10,1)=



1
Expert's answer
2021-09-16T00:46:24-0400
"C(6,1)=\dbinom{6}{1}=\dfrac{6!}{1!(6-1)!}=\dfrac{6}{1}=6"

"C(8,1)=\dbinom{8}{1}=\dfrac{8!}{1!(8-1)!}=\dfrac{8}{1}=8"

"C(10,1)=\dbinom{10}{1}=\dfrac{10!}{1!(10-1)!}=\dfrac{10}{1}=10"